Spindler, Inc. (a U.S.-based company), imports surfboards from a supplier in Brazil and sells them in the United States. Purchases are denominated in terms of the Brazilian real (BRL). During 2020, Spindler acquires 210 surfboards at a price of BRL 1,600 per surfboard, for a total of BRL 336,000.00. Spindler will pay for the surfboards when it sells them. Relevant exchange rates are as follows:

Date U.S. Dollar per
Brazilian Real (BRL)
September 1, 2020 $ 0.240
December 1, 2020 0.230
December 31, 2020 0.250
March 1, 2021 0.235
  1. Assume that Spindler acquired the surfboards on September 1, 2020, and made payment on December 1, 2020. What is the effect of the exchange rate fluctuations on reported income in 2020?
  2. Assume that Spindler acquired the surfboards on December 1, 2020, and made payment on March 1, 2021. What is the effect of the exchange rate fluctuations on reported income in 2020 and 2021?
  3. Assume that Spindler acquired the surfboards on September 1, 2020, and made payment on March 1, 2021. What is the effect of the exchange rate fluctuations on reported income in 2020 and in 2021?
Answer #1

a. Acquired surfboards on September 1, 2020 - 210 Surfboards @ BRL 1,600 = BRL 336,000.

Exchange rate was $0.240 / BRL on September 1, 2020 i.e. $ Payable = 336,000 * 0.240 = $80,640

Payment made on December 1, 2020 when exchange rate was $0.230 / BRL i.e.

Actual $ outflow = 336,000 * 0.230 = $77,280

Foreign Exchange rate gain of $ 3,360 ($80,640 - $77,280) will be reported in income of 2020.

b. Acquired surfboards on Decemberr 1, 2020 - 210 Surfboards @ BRL 1,600 = BRL 336,000.

Exchange rate was $0.230 / BRL on December 1, 2020 i.e. $ Payable = 336,000 * 0.230 = $77,280

Since on December 31, closing of Financial year, supplier payable balance should be adjusted in the financial statement and hence BRL 336,000 should be converted at a closing rate of $0.250 / BRL i.e $84,000 (336,000 * 0.250). Hence Unrealized loss on foreign exchange of $ 6,720 ($84,000 - $77,280) will be reported in income statement in 2020.

Payment made on March 1, 2021 when exchange rate was $0.235 / BRL i.e.

Actual $ outflow = 336,000 * 0.235 = $78,960

Foreign Exchange gain of $ 5,040 ($ 84,000 - $78,960) will be reported in income statement in 2021.

c. Acquired surfboards on September 1, 2020 - 210 Surfboards @ BRL 1,600 = BRL 336,000. .

Exchange rate was $0.240 / BRL on September 1, 2020 i.e. $ Payable = 336,000 * 0.240 = $80,640

Since on December 31, closing of Financial year, supplier payable balance should be adjusted in the financial statement and hence BRL 336,000 should be converted at a closing rate of $0.250 / BRL i.e $84,000 (336,000 * 0.250). Hence Unrealized loss on foreign exchange of $ 3,360 ($84,000 - $80,640) will be reported in income statement in 2020.

Payment made on March 1, 2021 when exchange rate was $0.235 / BRL i.e.

Actual $ outflow = 336,000 * 0.235 = $78,960

Foreign Exchange gain of $ 1,680 ($ 80,640 - $78,960) will be reported in income statement in 2021.